OpenStudy (jdfkljositjr):

Solve the triangle. A = 33°, a = 20, b = 13

OpenStudy (zpupster):

is it a right triangle

OpenStudy (zpupster):

 a^2 = b^2 = c^2 pyth.theorem (20)^2 + (13)^2 = c^c 400 + 169 = c^2 c= sqrt 569